OpenStudy (anonymous):

what is the sum o 4/5, 8/15 and 1/3

OpenStudy (unklerhaukus):

well the common denominator is 15, convert all the fractions to 15ths

OpenStudy (karatechopper):

ok well u have to denominators the same so change the 1/3 to a common denominator does 3 go into 15?

OpenStudy (anonymous):

25/15 or 1 and 10/15 simplified to 1 and 2/3

OpenStudy (anonymous):

15

OpenStudy (anonymous):

15

OpenStudy (karatechopper):

k so then 1/3=?/15

OpenStudy (unklerhaukus):

4/5, 8/15 and 1/3 =12/15, 8/15 and 5/15 =25/15 =5/3

OpenStudy (anonymous):

5/15

OpenStudy (karatechopper):

then u have the 4/5=?/15?

OpenStudy (anonymous):

12/15

OpenStudy (karatechopper):

k so now u 12/15+8/15+5/15=?

OpenStudy (anonymous):

25/15

OpenStudy (karatechopper):

good! now i gotta eat bbs!

OpenStudy (anonymous):

I have more